OpenStudy (anonymous):

I will fan and medal! Solve the following system and show all of your work. 3y − 2x = 11 y + 2x = 9

OpenStudy (anonymous):

Using the Elimination Method: x can be eliminated by adding the two equations 3y + y - 2x + 2x = 11 + 9 4y + 0x = 20 4y = 20 y = 20/4 = 5 Using y + 2x = 9 ; replace ( ie substitute) y = 5 5 + 2x = 9 2x = 9 - 5 = 4 x = 4/2 = 2 Using the Substitution Method y = 9 - 2x Replace (Substitute) y in the equation 3y - 2x = 11 with 9 - 2x 3(9 - 2x) - 2x = 11 27 - 6x - 2x = 11 27 - 8x = 11 - 8x = 11 - 27 - 8x = -16 x = -16/-8 x = 2 Replace x with 2 in the equation y = 9 - 2x to find for y y = 9 - 2(2) y = 9 - 4 y = 5

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

No Problem :)

OpenStudy (anonymous):

u can also substitute lolz :)